1. Which of the following is true?

A. Total liabilities is made up of current liabilities and long term debt.

B. EBT is EBIT minus depreciation and amortization.

C. Interest expense is paid to common shareholders.

D. Total assets is made up of current assets, preferred stock, and gross fixed assets.

2. Assume EBIT is 6, Net Income is 3, the flat tax rate is 40% and the company has no preferred stock. What is Earnings Before Taxes?

______________________

3. Assume EBIT is 4, Net Income is 2, the flat tax rate is 20% and the company has no preferred stock. What is the interest expense?

A. 1.50

B. 2.00

C. 3.25

D. 2.50

4. Assume Total Assets are 10 and New Fixed Assets are 7. What are the Current Assets?

A. 17.00

B. 1.43

C. 3.00

D. 0.70

Thank you in advance !!
Answer #1

1.

A. Total liabilities is made up of current liabilities and long term debt

2.

Net income equals earnings after tax, if it is 3 and tax rate is 40%, then earnings before tax is

=Earnings after tax / (1-tax rate)

=3 / 0.6

=5

3.

If net income is 2 ,and tax rate is 20%, earnings before tax is:

=2/0.8

=2.5

If EBIT is 4 and EBT is 2.5, interest component is:

=EBIT - EBT

=4 - 2.5

=1.5

4.

Total assets = Fixed assets + Current assets

10 = 7 + current assets

Current assets = 3